OpenStudy (anonymous):

at a $5000 investment earns $550 annual simple interest in one year.What is the annual interest rate? Please help me with this.The ossible answers are A.11% B.12% C.10.5 D.11.5

OpenStudy (sidsiddhartha):

use I=PRT

OpenStudy (sidsiddhartha):

550 = 5000*R*1 550/5000 = 5000r/5000 0.11 = r = 11%